OpenStudy (anonymous):

How would you explain to a student that 1*3*5*...(2n+1) = 1*3*5*...(2n-1)(2n+1)?

OpenStudy (anonymous):

Specifically, I have a few students struggling with understanding that \[\frac{1\cdot3\cdot5 \dotsm(2n+1)}{1\cdot3\cdot5\dotsm(2n-1)}=2n+1.\]

OpenStudy (zarkon):

I would write out more of the end terms... \[\frac{1\cdot3\cdot5 \dotsm(2n+1)}{1\cdot3\cdot5\dotsm(2n-1)}\] \[=\frac{1\cdot3\cdot5 \dotsm (2n-3)(2n-1)(2n+1)}{1\cdot3\cdot5\dotsm(2n-3)(2n-1)}\]

OpenStudy (trojanpoem):

Maybe assume 1.3.5...... until (2n - 3) = k \[\frac{ k (2n - 1)(2n+1) }{ k(2n-1) }\] (I am not a teacher)